POSITIVE DEPENDENCE ORDERINGS AND STOPPING TIMES

BRUNO BASSAN^{1} AND MARCO SCARSINI^{2}

1Dipartimento di Matematica, Politecnico di Milano, P.le Leonardo da Vinci 32, I-20133 Milano, Italy 2Dipartimento di Metodi Quantitativi, Università G. D’Annunzio, viale Pindaro 42, I-65127 Pescara, Italy

(Received November 9, 1992; revised August 9, 1993)

Abstract. We study the positive dependence of pairs of stochastic processes and examine its relation with the properties of certain stopping times. Some special cases, such as dependent random walks, Gaussian processes and exchangeable sequences of elliptically contoured random variables, are taken into account.

Key words and phrases: Stochastic orderings, bivariate random walks, bivariate Gaussian processes, exchangeable sequences.